OpenStudy (yrivers36):

Solve each equation symbolically. With explanation please. sqrt (x)+1=sqrt (x+2)

OpenStudy (anonymous):

hint... \[\LARGE \sqrt x+1=\sqrt{x+2}\] square both sides..

OpenStudy (yrivers36):

Im not sure but I think I should foil left side of problem

OpenStudy (yrivers36):

\[(\sqrt{x}+1)(\sqrt{x}+1)-x+2\]

OpenStudy (yrivers36):

thats an = in front of x

OpenStudy (anonymous):

\[\LARGE (\sqrt x+1)^2=(\sqrt{x+2})^2\] \[\LARGE (\sqrt x)^2+2\cdot \sqrt{x}\cdot 1+1^2=x+2\] \[\LARGE x+2\cdot \sqrt{x}=x+2-1\] \[\LARGE 2\cdot \sqrt{x}=x-x +2-1\] \[\LARGE 2\cdot \sqrt{x}=1\] \[\LARGE \sqrt{x}=\frac12\] square both sides and there you go :)

OpenStudy (anonymous):

can you do it ?

OpenStudy (yrivers36):

solution manuel says x = 1/4

OpenStudy (anonymous):

that's correct. Well done ! ;)

OpenStudy (yrivers36):

Thank you once again :)

OpenStudy (anonymous):

Glad to help :)